Random samples of size 4 are drawn with replacement form finite population 3, 6, 9, 20. What is the variance of the sample?

1
Expert's answer
2022-05-12T09:17:53-0400

Population mean:

"\mu=\cfrac{3+6+9+20}{4}=9.5."


Population variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\\
=\begin{Bmatrix}
 3-9.5,6-9.5,9-9.5,20-9.5
\end{Bmatrix}="

"=\begin{Bmatrix}
-6.5, - 3.5,-0.5,10.5
\end{Bmatrix},"

"\sigma^2=(-6.5)^2\cdot \cfrac{1}{4}+(-3.5)^2\cdot \cfrac{1}{4}+\\
+(-0.5)^2\cdot \cfrac{1}{4}+10.5^2\cdot \cfrac{1}{4}=41.25."


Variance of the sampling distribution of sample means:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{41.25}{4}=10.3125."